A strong spring is mounted to a steel I-beam so that the spring hangs vertically. A weight is hung from the spring. The energy P (arbitrary units) stored in the spring is given by P(x) = 2x^2, where x is the distance in inches that the end of the spring is stretched from its equilibrium position. Find P(4).
a. 4 units
b. 16 units
c. 8 units
d. 32 units

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(57424)

P(x) = 2x^2
P(4)=2(4^2)=32
Cheers,
Stan H.

Answer by checkley71(8403)

P(X)=2X^2 OR P(4)=2*4^2 OR P(4)=2*16 OR P(4)=32 UNITS.